# The Rise of Small Language Models in Healthcare: A Comprehensive Survey

### v1

@@ 3. Prompting, structured fine-tuning and reasoning: Adapting SLMs to clinical precision
After pretraining and the integration of attention mechanisms, prompt engineering, reasoning and instruction tuning becomes the most essential components to adapt the general instruct-language model to specific healthcare tasks. Fine-tuning enables the model to align with domain-specific objectives and improve performance in clinical settings. In this section, we discuss in-context learning (§3.1) and fine-tuning (§3.2).
@@ 3.1. In-context learning
In-context learning (ICL) in healthcare is the ability of language models to learn and adapt to new healthcare NLP tasks and provide meaningful outputs based on the context in clinical notes, without requiring explicit retraining or fine-tuning. While Dong et al. (2022) provides a broad overview of ICL, a comprehensive survey dedicated to its applications in cNLP remains absent. In-context learning enables models to adapt to tasks without gradient updates, while preference alignment further refines their behavior to align with human values and domain-specific expectations, ensuring safer and more relevant responses in healthcare.
Alignment: The most renowned method for preference alignment is reinforcement learning from human feedback (RLHF), which relies on reward models and policy updates that often introduce instability due to probability ratio-based optimization (Bai et al., 2022). DeepSeek’s Group Relative Policy Optimization (GRPO) made RLHF more efficient (Shao et al., 2024), especially for complex reasoning tasks, while Stanford’s direct preference optimization (DPO) refines clinician response selection by dynamically adjusting the log probabilities of preferred versus non-preferred responses. These models uses human feedback to train language models and generate more preferred responses without requiring a separate reward model. Past studies reported that supervised fine-tuning (SFT) is sufficient for healthcare CLS tasks, whereas direct performance optimization yields superior results for clinical reasoning, and clinical triage (Savage et al., 2024). Other than DPO (Ahn et al., 2024), the research community has explored multi-agent approach of preferential data annotation through multiple rounds of dialogue (Zhang et al., 2025a; Gururajan et al., 2024) and reinforcement learning with environmental feedback (Liu et al., 2024b) in healthcare applications. Thus, preferential alignment methods have the potential to alleviate the labor-intensive workload of clinicians.
@@
Reasoning: Reasoning further increases the capabilities of model to infer relationships between concepts, drawing conclusions, and applying domain knowledge to produce outputs that are logically consistent with medical standards. Huang and Chang (2022) discuss different mechanisms for improving and eliciting reasoning in LLMs and benchmarks methods for evaluating reasoning abilities. To encourage reasoning in LLMs, chain-of-thought (CoT) prompting (Wei et al., 2022) and “let us think step-by-step" prompting (Kojima et al., 2022) guides them to generate explicit reasoning steps before answering. CoT allows structured inference by breaking down complex medical queries into intermediate reasoning steps in healthcare NLP tasks (Wang et al., 2024e; Jiang et al., 2025). Wu et al. (2025a) critiques CoT prompting for its limitations across subspecialties and k-shot approaches for introducing noise, proposing AutoMedPrompt, leveraging textual gradients for optimizing system prompts to enhance medical reasoning in foundation LLMs.
@@
The integration of external knowledge has advanced healthcare informatics by enhancing factual accuracy and reducing hallucinations (Li et al., 2025; Wang et al., 2024c). Retrieval-augmented generation (RAG) further strengthens healthcare language models by incorporating clinical knowledge into text generation to deliver contextually relevant responses (Lahiri and Hu, 2024). The integration of RAG and CoT demonstrates improved performance on high-quality clinical notes, where early retrieval effectively anchors subsequent reasoning steps in domain-specific evidence while CoT-driven RAG offers advantages in handling lengthy and noisy notes by guiding retrieval through structured reasoning (Wu et al., 2025b; Li et al., 2024e). In MetaICL (2023), meta-learns task-agnostic prompts for generalizing across specialties (e.g., cardiology to nephrology), “learning" patterns dynamically from the context (Min et al., 2021). The integration of structured knowledge from a knowledge graph (KG) into the model’s training or inference process enhance the model’s ability to understand context and reason over relationships. KG has the potential to map patient data to structured ontologies (e.g., SNOMED-CT, UMLS) and infer relationships (e.g., "fatigue + jaundice → hepatitis risk"). Recent studies on neuro-symbolic AI suggest the need of further investigation of metacognition121212model’s capacity to monitor, assess, and adapt its own reasoning or output generation processes, often by estimating confidence, recognizing uncertainty, or reflecting on errors. for advancing the field towards more intelligent, reliable, and context-aware AI systems (Colelough and Regli, 2025; Didolkar et al., 2024). Metacognition improves decision-making in healthcare practitioners through cognitive bias identification, facilitation of systematic approaches, holistic approaches, and creative thinking (Church and Carroll, 2023). MD-PIE emulates cognitive and reasoning abilities in medical reasoning and decision-making (Esteitieh et al., 2025). However, more robust evaluation frameworks that incorporate metacognitive abilities are required as they are essential for developing reliable language model enhanced clinical decision support systems (Griot et al., 2025).
3.2. Supervised fine-tuning
@@
Fine-tuning on smaller, task-specific datasets enables language models to effectively leverage pre-trained knowledge for specialized domains or tasks (Fu et al., 2024; Tran et al., 2024). These techniques are address challenges like limited computing resources, data quality, availability, and robustness, ensuring efficient adaptation to new tasks without extensive retraining. Instruction tuning is sometimes used interchangeably with instruction fine-tuning; however, the latter often refers to training with chat-specific templates. While the distinction is debated in the literature, we do not focus on this differentiation in this work.
@@
Instruction tuning: Instructing tuning leverages instructions for flexibility. Instruction datasets are created by integrating annotated data into instruction-output pairs using templates. Alternatively, LLMs can generate outputs from manually or seed-generated instructions. In the human-generated questions with LLM annotations approach, humans create diverse, relevant questions, while LLMs provide initial answers. These answers are then reviewed and refined by humans, resulting in high-quality datasets for instruction tuning. The LLM-generated SDHG for instruction tuning increases training example diversity and reduces manual effort. Nazar et al. (2025) provides A detailed guidance on designing and evaluating instruction datasets for healthcare is available appropriate data construction methods and strategies ensuring data quality (Zhang et al., 2023a; Nazar et al., 2025). BioInstruct (Tran et al., 2024) and MIMIC-Instr (Wu et al., 2024) are the high-quality healthcare instruction datasets designed to enhance the training of language models for clinical NLP tasks.
@@
Instruction fine-tuning Instruction fine-tuning leverages focused instructions for specific tasks. A recent study suggests that while full parameter instruction fine-tuning can enhance task-specific performance, it may also lead to knowledge degradation and an increased risk of hallucination, potentially due to the overwriting of pre-trained knowledge during the fine-tuning process (Ghosh et al., 2024). AlpaCare is fine-tuned on MedInstruct-52k, a diverse machine-generated medical instruction-finetuning dataset, that outperforms previous medical LLMs (Zhang et al., 2023b). LlamaCare enhances LLMs’ ability to predict outcomes by fine-tuning existing LLM on various clinical tasks (Li et al., 2024c). Instruction fine-tuning is employed to adapt the SLM for biomedical tasks (Luo et al., 2024), medical entity recognition (Liu et al., 2024c), radiology tasks (Ranjit et al., 2024a), ICU readmissions (Namvar et al., 2025), medical documentation (Leong et al., 2024) and billing and coding (Rollman et al., 2025). However, the high-quality, diverse, and relevant instructions are key for SLM fine-tuning.
@@
Parameter efficient fine tuning (PEFT) adapts models to specific tasks by training only a small (upto <5%) subset of parameters making them deployable on edge devices such as tablets in rural clinics. This rapid customization fine-tunes the model more efficiently while preserving the advantages of the pre-trained model. Starting with LoRA, subsequent methods like LoRA+, LoRA-FA, and AdaLoRA improve efficiency, memory usage, and adaptability, while approaches such as LoRA-drop, and DoRA introduce novel strategies like random matrix adaptation, output-based pruning, and weight decomposition to further enhance parameter-efficient tuning (Mao et al., 2025). LoRA injects low-rank matrices into existing weights, allowing models to adapt to specialized tasks without full retraining, preserving foundational medical knowledge with minimal overhead (Shi et al., 2024). Recent fine-tuning of Phi3.5-mini [^13] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct through quantized low-rank adapters (QLoRA) underscores its potential for providing basic medical recommendations in telemedicine and chatbot systems. Adapter modules, inserted between transformer layers, facilitate multi-task learning on these small layers instead of modifying the base model as mixture of experts (Cai et al., 2025). Domain-specific PEFT approach (Gajulamandyam et al., 2025) is employed for hospital discharge paper SUM (Goswami et al., 2024), clinical RE (Zhou et al., 2023a), and biomedical NER (Abishek et al., 2024). Prompt tuning optimizes continuous “soft prompts" to guide models toward structured outputs (Wang et al., 2023c; He et al., 2024). In prefix tuning a small set of trainable parameters (prefixes) is prepended to the input representation in a frozen language model, allowing adaptation to domain-specific tasks without modifying the core model weights.
@@ 4. Compression techniques: Ensuring accessibility and sustainability
The need for compressed models—via quantization, pruning, and knowledge distillation—arises from the demand to deploy LLMs efficiently on resource-constrained devices, reduce inference latency and memory usage, cut energy costs, and enable real-time applications while maintaining performance close to their larger counterparts. In healthcare, compressed models enable real-time clinical decision support, preserving patient privacy while operating on limited hardware for NLP-driven clinical decision-making.
@@ 4.1. Knowledge distillation
Knowledge distillation (KD) is a model compression technique that trains a smaller student model to replicate the behavior of a larger teacher model (Hinton et al., 2015). KD has proven effective for creating high-performance SLMs from LLMs while significantly reducing model size (Wang et al., 2024d). KD is comprised of two components: knowledge and distillation. The knowledge covers a variety of teacher-student frameworks: labeling (teacher-generated outputs), expansion (in-context sample generation), data curation (meta-driven synthesis), feature extraction (teacher-derived internal representations), feedback (evaluative guidance on student outputs), and self-knowledge (student-led output generation and quality filtering) (Xu et al., 2024). Distillation covers supervised fine-tuning, divergence and similarity, Reinforcement learning, rank optimization. In supervised fine-tuning, also known as sequence-level KD, the student model is trained to maximize the likelihood of teacher-generated sequences, aligning its predictions with the teacher’s outputs through divergence or similarity-based loss functions. Similarity-based methods aligns internal representations of teacher and student models while divergence-based methods minimize distributional differences. Reinforcement learning-based distillation leverages teacher-generated feedback to train reward models and optimize student policies, aligning student outputs with teacher preferences through reward maximization and divergence minimization. Ranking optimization offers a stable, sample-efficient alternative to reinforcement learning by directly fine-tuning language models to favor preferred outputs over less favored ones using fixed teacher-generated preference datasets. Furthermore, the KD methods are categorized into white-box vs. black-box based on the information accessible from the teacher (Yang et al., 2024c).
@@
In white-box KD, a student not only observes the teacher’s outputs but also has access to the teacher model’s internal parameters or representations (Agarwal et al., 2024). This additional insight into the teacher’s “internal structure" often enables the student to perform better since it can directly leverage the teacher’s learned features and hidden knowledge, addressing the challenges of LLM scalability and portability for health event prediction (Ding et al., 2024). A feature-level knowledge distillation technique was introduced for medication recommendation, thereby training the SLM using hidden states of LLM (Liu et al., 2024d). In black-box KD, by contrast, only the teacher’s outputs (e.g. predicted logits or generated responses) are available to the student; the “internal structure" of the teacher is treated as a black box (Peng et al., 2023; Wang et al., 2023b). This scenario commonly arises when using proprietary LLM APIs as teacher models, where only input–output pairs are accessible, but the underlying model weights remain unavailable. Despite the limited information, black-box KD has shown promising results by fine-tuning students on the input–output behaviors of large models. Complex reasoning and problem-solving skills can be transferred by distilling the teacher’s CoT and prompt-based capabilities.
@@ 5.1. Current state-of-the-art
We assemble the evaluation of existing healthcare SLMs on various NLP tasks, comparing their performance in medical QA through accuracy in Fig. 4(a), NLU through F-score in Fig. 4(b), and mental health analysis through weighted F-measure in Fig. 4(c). We do not claim any computational novelty, but carry over the results from existing studies to report the current state-of-the-art.
@@
Medical question answering: Among all others, the results of six SLMs/versions of healthcare-related models developed for medical tasks (BioMedLM (Bolton et al., 2024), BioMistral-SLERP (Labrak et al., 2024), BioMistral-DARE (Labrak et al., 2024), MediTron (Chen et al., 2023), Med-$R^{2}$ (Lu et al., 2025), PMC-LLAMA (Wu et al., [n. d.])) were comparable on three most commonly used medical QA datasets: PubMedQA (Jin et al., 2019), MedMCQA (Pal et al., 2022), and MedQA-USMLE (Kung et al., 2023). In addition to the medical tasks, we collected results from conversational models in healthcare (ChatDoctor (Li et al., 2023a), MedAlpaca (Han et al., 2023), and AlpaCare (Zhang et al., 2023b)) Fig. 4(a) shows no single model dominates across all medical QA datasets. PMC-LLAMA achieves the highest score on PubMedQA (77.09%), demonstrating strong factual recall in biomedical literature. On MedQA-USMLE, which emphasizes clinical reasoning, Med-$R^{2}$ significantly outperforms all others with a score of 81.06%, indicating its robustness in handling complex medical exam-style questions. BioMedLM shows the most consistent performance across all datasets, ranking among the top models for MedMCQA (57.3%), a multiple-choice QA dataset focused on general medical knowledge. These results underscore the importance of task-specific fine-tuning and diverse pretraining corpora in building high-performing medical language models. The effectiveness of healthcare SLMs varies across datasets, reflecting differences in reasoning complexity and domain specificity. The lack of a universally dominant model highlights the need for adaptive SLM architectures for clinical settings. Incorporating modular training strategies or task-specific adapters could help bridge the performance gap across diverse datasets.
@@
Mental health analysis: We gather the results from three models: MentaLLAMA (Yang et al., 2024b), mhGPT (Kim et al., 2024a), and MentalQLM (Shi et al., 2024); among which MentaLLAMA have two variants: base model and chat model. The more widely adapted datasets are interpersonal risk factors (Garg et al., 2023), Dreddit (Turcan and McKeown, 2019), SAD (Mauriello et al., 2021), MultiWD (Garg et al., 2024), and DR (Pirina and Çöltekin, 2018). MentalQLM achieves the highest scores on IRF (76.81%), Dreddit (82.85%), SAD (67.06%), and MultiWD (77.88%), reflecting its robustness in identifying psychological distress, sentiment, and well-being markers as shown in Fig. 4(c). For the DR dataset, which lacks a score for mhGPT, MentalQLM narrowly outperforms other models with a top score of 84.20%. MentaLLAMA-chat-7B also shows competitive performance, particularly in DR (83.95%) and MultiWD (75.79%), indicating its strength in conversational contexts. The superior performance by existing SLMs underscores the importance of contextual and sensitive language understanding in mental health applications. ChatCounselor is benchmarked against existing LLMs using a robust GPT-4–based evaluation protocol, which emphasizes natural, human-like dialogue over mechanistic response listing (Liu et al., 2023a). ChatCounselor outperforms most open-source models and approaches ChatGPT in quality, underscoring the impact of domain-specific fine-tuning on enhancing personalized mental health support. Future work should explore benchmarking clinical mental health SLMs against existing mental health–focused or general-purpose LLMs to strengthen the confidence of key stakeholders- healthcare professionals. Moreover, as the training datasets are primarily derived from social media, the development of a well-established model is essential for effective deployment in clinical or hospital settings.
Miscellaneous: We further observed three biomedical language models—BioMistral, BioMistral-NLU, and BioGPT—on the RE task using the F1 score as the evaluation metric. BioGPT outperforms the other healthcare SLMs. The ClinicalMamba 2.8B achieves an F1 score of 56.51 and 74.34 for IE task task of rare and common ICD coding, respectively. The performance of various language models on NER tasks across six biomedical datasets demonstrates superior performance by BioMistral-NLU. However, the F1 score varies from 55% to 90% for different NER dataset, highlight the need of specialized medical NER models. Clinical conditions (e.g., cancer) and medical expertise (e.g., radiology) consider the construction of specialized models using diagnosis generation tasks/RE, and radiology NLI/IE, respectively. The current state of the art highlights the need for language models that achieve an optimal balance between computational efficiency and high predictive accuracy, while ensuring responsible use of resources.

### v2

@@ 3. Prompting, structured fine-tuning and reasoning: Adapting SLMs to clinical precision
After pretraining and the integration of attention mechanisms, prompt engineering, reasoning and instruction tuning becomes the most essential components to adapt the general instruct-language model to specific healthcare tasks. Fine-tuning enables the model to align with domain-specific objectives and improve performance in hospital settings. In this section, we discuss in-context learning (§3.1) and fine-tuning (§3.2).
@@ 3.1. In-context learning
In-context learning (ICL) in healthcare is the ability of language models to learn and adapt to new healthcare NLP tasks and provide meaningful outputs based on the context in clinical notes, without requiring explicit retraining or fine-tuning. While Dong et al. (2022) provide a broad overview of ICL, a comprehensive survey dedicated to its applications in cNLP remains absent. In-context learning enables models to adapt to tasks without gradient updates, while preference alignment further refines their behavior to align with human values and domain-specific expectations, ensuring safer and more relevant responses in healthcare.
Alignment: The most renowned method for preference alignment is reinforcement learning from human feedback (RLHF), which relies on reward models and policy updates that often introduce instability due to probability ratio-based optimization (Bai et al., 2022). DeepSeek’s Group Relative Policy Optimization (GRPO) made RLHF more efficient (Shao et al., 2024), especially for complex reasoning tasks, while Stanford’s direct preference optimization (DPO) refines clinician response selection by dynamically adjusting the log probabilities of preferred versus non-preferred responses. These models uses human feedback to train language models and generate more preferred responses without requiring a separate reward model. Past studies reported that supervised fine-tuning (SFT) (§3.2) is sufficient for healthcare CLS tasks, whereas direct performance optimization yields superior results for clinical reasoning, and clinical triage (Savage et al., 2024). Other than DPO (Ahn et al., 2024), the research community has explored multi-agent approach of preferential data annotation through multiple rounds of dialogue (Zhang et al., 2025a; Gururajan et al., 2024) and reinforcement learning with environmental feedback (Liu et al., 2024b) in healthcare applications. Thus, preferential alignment methods have the potential to alleviate the labor-intensive workload of clinicians.
@@
Reasoning: Reasoning further increases the capabilities of model to infer relationships between concepts, drawing conclusions, and applying domain knowledge to produce outputs that are logically consistent with medical standards. Huang and Chang (2022) discussed different mechanisms for improving and eliciting reasoning in LLMs and benchmarks methods for evaluating reasoning abilities. To encourage reasoning in LLMs, chain-of-thought (CoT) prompting (Wei et al., 2022) and “let us think step-by-step” prompting (Kojima et al., 2022) guides them to generate explicit reasoning steps before answering. CoT allows structured inference by breaking down complex medical queries into intermediate reasoning steps in healthcare NLP tasks (Wang et al., 2024e; Jiang et al., 2025). Wu et al. (2025a) critiques CoT prompting for its limitations across subspecialties and k-shot approaches for introducing noise, proposing AutoMedPrompt, leveraging textual gradients for optimizing system prompts to enhance medical reasoning in foundation LLMs.
@@
The integration of external knowledge has advanced healthcare informatics by enhancing factual accuracy and reducing hallucinations (Li et al., 2025; Wang et al., 2024c). Retrieval-augmented generation (RAG) further strengthens healthcare language models by incorporating clinical knowledge into text generation to deliver contextually relevant responses (Lahiri and Hu, 2024). The integration of RAG and CoT demonstrates improved performance on high-quality clinical notes, where early retrieval effectively anchors subsequent reasoning steps in domain-specific evidence while CoT-driven RAG offers advantages in handling lengthy and noisy notes by guiding retrieval through structured reasoning (Wu et al., 2025b; Li et al., 2024e). In MetaICL, meta-learns task-agnostic prompts for generalizing across specialties (e.g., cardiology to nephrology), “learning” patterns dynamically from the context (Min et al., 2021). The integration of structured knowledge from a knowledge graph (KG) into the model’s training or inference process enhance the model’s ability to understand context and reason over relationships. KG has the potential to map patient data to structured ontologies (e.g., SNOMED-CT) and infer relationships (e.g., "fatigue + jaundice → hepatitis risk"). Recent studies on neuro-symbolic AI suggest the need of further investigation of metacognition121212model’s capacity to monitor, assess, and adapt its own reasoning or output generation processes, often by estimating confidence, recognizing uncertainty, or reflecting on errors. for advancing the field towards more intelligent, reliable, and context-aware AI systems (Colelough and Regli, 2025; Didolkar et al., 2024). Metacognition improves decision-making in healthcare practitioners through cognitive bias identification, facilitation of systematic approaches, holistic approaches, and creative thinking (Church and Carroll, 2023). MD-PIE emulates cognitive and reasoning abilities in medical reasoning and decision-making (Esteitieh et al., 2025). However, more robust evaluation frameworks that incorporate metacognitive abilities are required as they are essential for developing reliable language model enhanced CDSS (Griot et al., 2025).
3.2. Supervised fine-tuning (SFT)
@@
Fine-tuning on smaller, task-specific datasets enables language models to effectively leverage pre-trained knowledge for specialized domains or tasks (Fu et al., 2024; Tran et al., 2024). These techniques address challenges like limited computing resources, data quality, availability, and robustness, ensuring efficient adaptation to new tasks without extensive retraining. Instruction tuning is sometimes used interchangeably with instruction fine-tuning; however, the latter often refers to training with chat-specific templates. While the distinction is debated in the literature, we do not focus on this differentiation in this work.
@@
Instruction tuning leverages instructions for flexibility. Instruction datasets are created by integrating annotated data into instruction-output pairs using templates. Alternatively, LLMs can generate outputs from manually or seed-generated instructions. In the human-generated questions with LLM annotations approach, humans create diverse, relevant questions, while LLMs provide initial answers. These answers are then reviewed and refined by humans, resulting in high-quality datasets for instruction tuning. The LLM-generated SDHG for instruction tuning increases training example diversity and reduces manual effort. Nazar et al. (2025) provide a detailed guidance on designing and evaluating instruction datasets for healthcare is available appropriate data construction methods and strategies ensuring data quality (Zhang et al., 2023a; Nazar et al., 2025). BioInstruct (Tran et al., 2024) and MIMIC-Instr (Wu et al., 2024) are the high-quality healthcare instruction datasets designed to enhance the training of language models for clinical NLP tasks.
@@
Instruction fine-tuning leverages focused instructions for specific tasks. A recent study suggests that while full parameter instruction fine-tuning can enhance task-specific performance, it may also lead to knowledge degradation and an increased risk of hallucination, potentially due to the overwriting of pre-trained knowledge during the fine-tuning process (Ghosh et al., 2024). AlpaCare is fine-tuned on MedInstruct-52k, a diverse machine-generated medical instruction fine-tuning dataset, that outperforms previous medical LLMs (Zhang et al., 2023b). LlamaCare enhances LLMs’ ability to predict outcomes by fine-tuning existing LLM on various clinical tasks (Li et al., 2024c). Instruction fine-tuning is employed to adapt the SLM for biomedical tasks (Luo et al., 2024), medical entity recognition (Liu et al., 2024c), radiology tasks (Ranjit et al., 2024a), ICU readmissions (Namvar et al., 2025), medical documentation (Leong et al., 2024) and billing and coding (Rollman et al., 2025). However, the high-quality, diverse, and relevant instructions are key for SLM fine-tuning.
@@
Parameter efficient fine tuning (PEFT) adapts models to specific tasks by training only a small (upto <5%) subset of parameters making them deployable on edge devices such as tablets in rural clinics. This rapid customization fine-tunes the model more efficiently while preserving the advantages of the pre-trained model. Starting with LoRA, subsequent methods like LoRA+, LoRA-FA, and AdaLoRA improve efficiency, memory usage, and adaptability, while approaches such as LoRA-drop, and DoRA introduce novel strategies like random matrix adaptation, output-based pruning, and weight decomposition to further enhance parameter-efficient tuning (Mao et al., 2025). LoRA injects low-rank matrices into existing weights, allowing models to adapt to specialized tasks without full retraining, preserving foundational medical knowledge with minimal overhead (Shi et al., 2024). Recent fine-tuning of Phi3.5-mini [^13] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct through quantized low-rank adapters (QLoRA) underscores its potential for providing basic medical recommendations in telemedicine and chatbot systems. Adapter modules, inserted between transformer layers, facilitate multi-task learning on these small layers instead of modifying the base model as mixture of experts (Cai et al., 2025). Domain-specific PEFT approach (Gajulamandyam et al., 2025) is employed for hospital discharge paper SUM (Goswami et al., 2024), clinical RE (Zhou et al., 2023a), and biomedical NER (Abishek et al., 2024). Prompt tuning optimizes continuous “soft prompts” to guide models toward structured outputs (Wang et al., 2023c; He et al., 2024). In prefix tuning a small set of trainable parameters (prefixes) is prepended to the input representation in a frozen language model, allowing adaptation to domain-specific tasks without modifying the core model weights.
@@ 4. Compression techniques: Ensuring accessibility and sustainability
The need for compressed models—via quantization, pruning, and knowledge distillation—arises from the demand to deploy LLMs efficiently on resource-constrained devices, reduce inference latency and memory usage, cut energy costs, and enable real-time applications while maintaining performance close to their larger counterparts. In healthcare, compressed models enable real-time CDSS, preserving patient privacy while operating on limited hardware for NLP-driven clinical decision-making.
@@ 4.1. Knowledge distillation
Knowledge distillation (KD) is a model compression technique that trains a smaller student model to replicate the behavior of a larger teacher model (Hinton et al., 2015). KD has proven effective for creating high-performance SLMs from LLMs while significantly reducing model size (Wang et al., 2024d). KD is comprised of two components: knowledge and distillation. The knowledge covers a variety of teacher-student frameworks: labeling (teacher-generated outputs), expansion (in-context sample generation), data curation (meta-driven synthesis), feature extraction (teacher-derived internal representations), feedback (evaluative guidance on student outputs), and self-knowledge (student-led output generation and quality filtering) (Xu et al., 2024). Distillation covers SFT, divergence and similarity, Reinforcement learning, rank optimization. In SFT, also known as sequence-level KD, the student model is trained to maximize the likelihood of teacher-generated sequences, aligning its predictions with the teacher’s outputs through divergence or similarity-based loss functions. Similarity-based methods aligns internal representations of teacher and student models while divergence-based methods minimize distributional differences. Reinforcement learning-based distillation leverages teacher-generated feedback to train reward models and optimize student policies, aligning student outputs with teacher preferences through reward maximization and divergence minimization. Ranking optimization offers a stable, sample-efficient alternative to reinforcement learning by directly fine-tuning language models to favor preferred outputs over less favored ones using fixed teacher-generated preference datasets. Furthermore, the KD methods are categorized into white-box vs. black-box based on the information accessible from the teacher (Yang et al., 2024c).
@@
In white-box KD, a student not only observes the teacher’s outputs but also has access to the teacher model’s internal parameters or representations (Agarwal et al., 2024). This additional insight into the teacher’s “internal structure” often enables the student to perform better since it can directly leverage the teacher’s learned features and hidden knowledge, addressing the challenges of LLM scalability and portability for health event prediction (Ding et al., 2024). A feature-level knowledge distillation technique was introduced for medication recommendation, thereby training the SLM using hidden states of LLM (Liu et al., 2024d). In black-box KD, by contrast, only the teacher’s outputs (e.g. predicted logits or generated responses) are available to the student; the “internal structure” of the teacher is treated as a black box (Peng et al., 2023; Wang et al., 2023b). This scenario commonly arises when using proprietary LLM APIs as teacher models, where only input–output pairs are accessible, but the underlying model weights remain unavailable. Despite the limited information, black-box KD has shown promising results by fine-tuning students on the input–output behaviors of large models. Complex reasoning and problem-solving skills can be transferred by distilling the teacher’s CoT and prompt-based capabilities.
@@ 5.1. Current state-of-the-art
We assemble the evaluation of existing healthcare SLMs on various NLP tasks, comparing their performance in medical QA through accuracy in Fig. 4(a), NLU through F1-score in Fig. 4(b), and mental health analysis through weighted F-measure in Fig. 4(c). We do not claim any computational novelty, but carry over the results from existing studies to report the current state-of-the-art.
@@
Medical question answering: Among all others, the results of six SLMs/versions of healthcare-related models developed for medical tasks (BioMedLM (Bolton et al., 2024), BioMistral-SLERP (Labrak et al., 2024), BioMistral-DARE (Labrak et al., 2024), MediTron (Chen et al., 2023), Med-$R^{2}$ (Lu et al., 2025), PMC-LLAMA (Wu et al., [n. d.])) were comparable on three most commonly used medical QA datasets: PubMedQA (Jin et al., 2019), MedMCQA (Pal et al., 2022), and MedQA-USMLE (Kung et al., 2023). In addition to the medical tasks, we collected results from conversational models in healthcare (ChatDoctor (Li et al., 2023a), MedAlpaca (Han et al., 2023), and AlpaCare (Zhang et al., 2023b)). Fig. 4(a) shows no single model dominates across all medical QA datasets. PMC-LLAMA achieves the highest score on PubMedQA (77.09%), demonstrating strong factual recall in biomedical literature. On MedQA-USMLE, which emphasizes clinical reasoning, Med-$R^{2}$ significantly outperforms all others with a score of 81.06%, indicating its robustness in handling complex medical exam-style questions. BioMedLM shows the most consistent performance across all datasets, ranking among the top models for MedMCQA (57.3%), a multiple-choice QA dataset focused on general medical knowledge. These results underscore the importance of task-specific fine-tuning and diverse pretraining corpora in building high-performing medical language models. The effectiveness of healthcare SLMs varies across datasets, reflecting differences in reasoning complexity and domain specificity. The lack of a universally dominant model highlights the need for adaptive SLM architectures for clinical settings. Incorporating modular training strategies or task-specific adapters could help bridge the performance gap across diverse datasets.
@@
Mental health analysis: We gather the results from three models: MentaLLAMA (Yang et al., 2024b), mhGPT (Kim et al., 2024a), and MentalQLM (Shi et al., 2024); among which MentaLLAMA have two variants: base model and chat model. The more widely adapted datasets are interpersonal risk factors (Garg et al., 2023), Dreddit (Turcan and McKeown, 2019), SAD (Mauriello et al., 2021), MultiWD (Garg et al., 2024), and Depression reddit (DR) (Pirina and Çöltekin, 2018). MentalQLM achieves the highest scores on IRF (76.81%), Dreddit (82.85%), SAD (67.06%), and MultiWD (77.88%), reflecting its robustness in identifying psychological distress, sentiment, and well-being markers as shown in Fig. 4(c). For the DR dataset, which lacks a score for mhGPT, MentalQLM narrowly outperforms other models with a top score of 84.20%. MentaLLAMA-chat-7B also shows competitive performance, particularly in DR (83.95%) and MultiWD (75.79%), indicating its strength in conversational contexts. The superior performance by existing SLMs underscores the importance of contextual and sensitive language understanding in mental health applications. ChatCounselor is benchmarked against existing LLMs using a robust GPT-4–based evaluation protocol, which emphasizes natural, human-like dialogue over mechanistic response listing (Liu et al., 2023a). ChatCounselor outperforms most open-source models and approaches ChatGPT in quality, underscoring the impact of domain-specific fine-tuning on enhancing personalized mental health support. Future work should explore benchmarking clinical mental health SLMs against existing mental health–focused or general-purpose LLMs to strengthen the confidence of key stakeholders- healthcare professionals. Moreover, as the training datasets are primarily derived from social media, the development of a clinical model is essential for effective deployment in hospital settings.
Miscellaneous: We further observed three biomedical language models—BioMistral, BioMistral-NLU, and BioGPT—on the RE task using the F1 score as the evaluation metric. BioGPT outperforms the other healthcare SLMs. The ClinicalMamba 2.8B achieves an F1 score of 56.51 and 74.34 for IE task task of rare and common ICD coding, respectively. The performance of various language models on NER tasks across six biomedical datasets demonstrates superior performance by BioMistral-NLU. However, the F1 score varies from 55% to 90% for different NER dataset, highlight the need of specialized medical NER models. Clinical conditions (e.g., cancer) and medical expertise (e.g., radiology) consider the construction of specialized models for diagnosis generation tasks/RE, and radiology NLI/IE, respectively. The current state of the art highlights the need for language models that achieve an optimal balance between computational efficiency and high predictive accuracy, while ensuring responsible use of resources.